OpenStudy (anonymous):

In five years, Josh will be 21 years old. How old is he now? Let J = Josh's current age. Which formula would calculate Josh's current age?

OpenStudy (lalaly):

J+5=21

OpenStudy (anonymous):

21-5=J

OpenStudy (anonymous):

21-5=j , or J+5=21